# Provably Efficient Multi-Objective Bandit Algorithms under Preference-Centric Customization

### v1

@@ D.6 Proof of Lemma 15
Lemma 18 (Eigenvalues of Sums of Hermitian Matrices [28], Eq.(11)).
@@
$\mathbb{E}[\boldsymbol{V}_{t}]$ $=\sum_{\ell=1}^{t}\mathbb{E}[w_{\ell}\boldsymbol{r}_{a_{\ell},\ell}\boldsymbol{r}_{a_{\ell},\ell}^{\top}]+\lambda\boldsymbol{I}=\sum_{i=1}^{K}\mathbb{E}\left[\sum_{\ell\in\mathcal{T}_{i,t}}w_{\ell}\boldsymbol{r}_{i,\ell}\boldsymbol{r}_{i,\ell}^{\top}\right]+\lambda\boldsymbol{I},$ (53)
@@
$\qquad=\boldsymbol{x}^{\top}\left(\tilde{w}\sum_{i=1}^{K}N_{i,t+1}\left(\boldsymbol{\mu}_{i}\boldsymbol{\mu}_{i}^{T}+\Sigma_{\boldsymbol{r,i}}\right)+\lambda\boldsymbol{I}-\sum_{i=1}^{K}\omega\sqrt{N_{i,t+1}\log\left(\frac{t}{\alpha}\right)}\boldsymbol{ee}^{\top}\right)\boldsymbol{x},$
@@
$\chi$ $\geq\boldsymbol{x}^{\top}\left(\underbrace{\tilde{w}\sum_{i=1}^{K}N_{i,t+1}\left(\boldsymbol{\mu}_{i}\boldsymbol{\mu}_{i}^{T}+\Sigma_{\boldsymbol{r,i}}\right)+\lambda\boldsymbol{I}}_{\mathbb{E}[\boldsymbol{V}_{t}]}-\sum_{i=1}^{K}\omega\sqrt{N_{i,t+1}\log\left(\frac{t}{\alpha}\right)}\boldsymbol{ee}^{\top}\right)\boldsymbol{x}$ (58)
=$\underset{(a)}{=}\boldsymbol{x}^{\top}\left(\tilde{w}t\boldsymbol{\tilde{\mu}}\boldsymbol{\tilde{\mu}}^{\top}+\left(\tilde{w}t\tilde{\sigma}^{2}_{r,t}+\lambda\right)\boldsymbol{I}-\sum_{i=1}^{K}\omega\sqrt{N_{i,t+1}\log\left(\frac{t}{\alpha}\right)}\boldsymbol{ee}^{\top}\right)\boldsymbol{x}$
@@
where (a) holds by Eq. 56 and Eq. 57, (b) holds since the squared root term is maximized when $N_{i,t+1}=t/K,\forall i\in[K]$. Note that $\boldsymbol{B}_{t}$ is diagonal matrix, and $-\boldsymbol{C}_{t}$ is rank-1 matrix yields one eigenvalue of $-\omega\sqrt{Kt\log\left(\frac{t}{\alpha}\right)}\|\boldsymbol{e}\|_{2}^{2}=-D\omega\sqrt{Kt\log\left(\frac{t}{\alpha}\right)}$ and $D-1$ eigenvalues of 0, we have
@@
Note that $\boldsymbol{A}_{t}$ and $-\boldsymbol{C}_{t}$ are rank-1 matrices and $\boldsymbol{B}_{t}$ is diagonal matrix, we can trivially derive that:
  • $\boldsymbol{A}_{t}+\boldsymbol{B}_{t}$ has one eigenvalue of $\tilde{w}t(\|\boldsymbol{\tilde{\mu}}\|_{2}^{2}+\tilde{\sigma}^{2}_{r,t})+\lambda$ and $D-1$ eigenvalues of $\tilde{w}t\tilde{\sigma}^{2}_{r,t}+\lambda$.
  • $-\boldsymbol{C}_{t}$ has one eigenvalue of $-\omega\sqrt{Kt\log\left(\frac{t}{\alpha}\right)}\|\boldsymbol{e}\|_{2}^{2}=-D\omega\sqrt{Kt\log\left(\frac{t}{\alpha}\right)}$ and $D-1$ eigenvalues of 0.
Since $\boldsymbol{A}_{t}+\boldsymbol{B}_{t}$ and $-\boldsymbol{C}_{t}$ are both symmetric, by applying Lemma 18, we have
@@ D.7 Proof of Lemma 17
$\sum_{t=M+1}^{T}\sqrt{\frac{\log\left(\frac{t}{\alpha}\right)}{N_{a_{t},t}}}$ $\underset{(a)}{\leq}\sqrt{\log\left(\frac{T}{\alpha}\right)}\sum_{i=1}^{K}\sum_{n=N_{i,M+1}}^{N_{i,T}}\sqrt{\frac{1}{n}}$ (62)
@@ D.8 Proof of Lemma 16
To derive the upper-bound of term $\sum_{t=M+1}^{T}\min\left(\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}\left[\boldsymbol{V}_{t-1}\right]^{-1}},\frac{1}{2\sqrt{2}}\right)$, we follow the similar techniques for analyzing the sum of instantaneous regret in OFUL [19]. Specifically, we first show that the sum of squared terms $\min\left(\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}\left[\boldsymbol{V}_{t-1}\right]^{-1}},\frac{1}{2\sqrt{2}}\right)^{2}$ is optimal up to $\mathcal{O}(\log T)$, and then extend the result to the sum of $\min\left(\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}\left[\boldsymbol{V}_{t-1}\right]^{-1}},\frac{1}{2\sqrt{2}}\right)$ using Cauchy-Schwarz inequality.
@@
$\emph{det}\left(\mathbb{E}[\boldsymbol{V}_{T}]\right)\geq\emph{det}\left(\mathbb{E}[\boldsymbol{V}_{M}]\right)\prod_{t=M+1}^{T}\left(1+\frac{\emph{det}\left(\Sigma_{r,a_{t}}\right)}{\emph{det}\left(\mathbb{E}[\boldsymbol{V}_{t-1}]\right)}+\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t-1}]^{-1}\boldsymbol{\mu}_{a_{t}}\right).$
Please see Appendix D.8.1 for the detailed proof of Lemma 20.
@@
Please see Appendix D.8.2 for the detailed proof of Lemma 21.
@@
$\sum_{t=M+1}^{T}\min\left(\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}\left[\boldsymbol{V}_{t-1}\right]^{-1}},\frac{1}{2\sqrt{2}}\right)^{2}$ (63)
$\qquad\qquad=\sum_{t=M+1}^{T}\min\left(\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t}]^{-1}\boldsymbol{\mu}_{a_{t}},\frac{1}{8}\right)$
$\qquad\qquad\underset{(a)}{\leq}\sum_{t=M+1}^{T}\frac{1}{8\log(9/8)}\log\left(1+\min\big{(}\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t}]^{-1}\boldsymbol{\mu}_{a_{t}},\frac{1}{8}\big{)}\right)$
$\qquad\qquad\leq\sum_{t=M+1}^{T}\frac{1}{8\log(9/8)}\log\left(1+\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t}]^{-1}\boldsymbol{\mu}_{a_{t}}\right)$
@@
$\log\left(\frac{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{T}]\right)}{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{M}]\right)}\right)\geq\sum_{t=M+1}^{T}\log\left(1+\frac{\text{det}\left(\Sigma_{r,a_{t}}\right)}{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{t-1}]\right)}+\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t-1}]^{-1}\boldsymbol{\mu}_{a_{t}}\right).$ (64)
@@
$D\log\left(1+\frac{\omega}{\lambda}(T-M)\right)\geq\log\left(\frac{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{T}]\right)}{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{M}]\right)}\right)\geq\sum_{t=M+1}^{T}\log\left(1+\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t}]^{-1}\boldsymbol{\mu}_{a_{t}}\right).$ (65)
@@
$\sum_{t=M+1}^{T}\min\left(\sqrt{\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t}]^{-1}\boldsymbol{\mu}_{a_{t}}},\frac{1}{2\sqrt{2}}\right)^{2}$ $\leq\frac{D}{8\log(9/8)}\log\left(1+\frac{\omega}{\lambda}(T-M)\right).$ (66)
@@
$\sum_{t=M+1}^{T}\min\Big{(}\sqrt{\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t}]^{-1}\boldsymbol{\mu}_{a_{t}}},\frac{1}{2\sqrt{2}}\Big{)}$ $\leq\sqrt{(T-M)\sum_{t=M+1}^{T}\min\Big{(}\sqrt{\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t}]^{-1}\boldsymbol{\mu}_{a_{t}}},\frac{1}{2\sqrt{2}}\Big{)}^{2}}$ (67)
@@ D.8.2 Proof of Lemma 21
For the proof of this lemma, we follow the main idea of Determinant-Trace Inequality in OFUL [19] (Lemma 10). Specifically, by the definition of $\boldsymbol{V}_{t}$, we have
$\log\left(\frac{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{T}]\right)}{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{M}]\right)}\right)$ $=\log\left(\text{det}\left(\frac{\mathbb{E}[\boldsymbol{V}_{M}]+\sum_{t=M+1}^{T}\mathbb{E}[w_{t}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}]}{\mathbb{E}[\boldsymbol{V}_{M}]}\right)\right)$ (72)
$\underset{(a)}{\leq}\log\left(\text{det}\left(1+\frac{\sum_{t=M}^{T}\mathbb{E}[w_{t}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}]}{\lambda\boldsymbol{I}}\right)\right)$
$=\log\left(\text{det}\left(1+\frac{1}{\lambda}\sum_{t=M+1}^{T}\mathbb{E}[w_{t}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}]\right)\right),$
where (a) holds since $\text{det}(\mathbb{E}[\boldsymbol{V}_{M}])\geq\text{det}(\mathbb{E}[\boldsymbol{V}_{0}])=\lambda\boldsymbol{I}$. Let $\xi_{1},...,\xi_{D}$ denote the eigenvalues of $\sum_{t=M+1}^{T}\mathbb{E}[w_{t}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}]$, and note:
$\sum_{d=1}^{D}\xi_{d}$ $=\text{Trace}\left(\sum_{t=M+1}^{T}\mathbb{E}[w_{t}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}]\right)$ (73)
$=\sum_{t=M+1}^{T}\mathbb{E}\left[\text{Trace}\left(w_{t}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}\right)\right]$
$=\sum_{t=M+1}^{T}\mathbb{E}\left[\text{Trace}\left(\frac{\omega}{\|\boldsymbol{r}_{a_{t},t}\|_{2}^{2}}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}\right)\right]$
$=\sum_{t=M+1}^{T}\mathbb{E}\left[\frac{\omega\|\boldsymbol{r}_{a_{t},t}\|_{2}^{2}}{\|\boldsymbol{r}_{a_{t},t}\|_{2}^{2}}\right]$
@@
$\log\left(\frac{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{T}]\right)}{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{M}]\right)}\right)$ $\leq\log\left(\text{det}\left(1+\frac{1}{\lambda}\sum_{t=M+1}^{T}\mathbb{E}[w_{t}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}]\right)\right)$

### v2

@@ D.6 Proof of Lemma 8
Lemma 11 (Eigenvalues of Sums of Hermitian Matrices (Fulton 2000), Eq.(11)).
@@
$\mathbb{E}[\boldsymbol{V}_{t}]$ $=\sum_{\ell=1}^{\top}\mathbb{E}[w_{\ell}\boldsymbol{r}_{a_{\ell},\ell}\boldsymbol{r}_{a_{\ell},\ell}^{\top}]+\lambda\boldsymbol{I}=\sum_{i=1}^{K}\mathbb{E}\left[\sum_{\ell\in\mathcal{T}_{i,t}}w_{\ell}\boldsymbol{r}_{i,\ell}\boldsymbol{r}_{i,\ell}^{\top}\right]+\lambda\boldsymbol{I},$ (30)
@@
$\qquad=\boldsymbol{x}^{\top}\left(\tilde{w}\sum_{i=1}^{K}N_{i,t+1}\left(\boldsymbol{\mu}_{i}\boldsymbol{\mu}_{i}^{\top}+\Sigma_{\boldsymbol{r,i}}\right)+\lambda\boldsymbol{I}-\sum_{i=1}^{K}\omega\sqrt{N_{i,t+1}\log\left(\frac{t}{\alpha}\right)}\boldsymbol{ee}^{\top}\right)\boldsymbol{x},$
@@
$\chi$ $\geq\boldsymbol{x}^{\top}\left(\underbrace{\tilde{w}\sum_{i=1}^{K}N_{i,t+1}\left(\boldsymbol{\mu}_{i}\boldsymbol{\mu}_{i}^{\top}+\Sigma_{\boldsymbol{r,i}}\right)+\lambda\boldsymbol{I}}_{\mathbb{E}[\boldsymbol{V}_{t}]}-\sum_{i=1}^{K}\omega\sqrt{N_{i,t+1}\log\left(\frac{t}{\alpha}\right)}\boldsymbol{ee}^{\top}\right)\boldsymbol{x}$ (35)
$\underset{(a)}{=}\boldsymbol{x}^{\top}\left(\tilde{w}t\boldsymbol{\tilde{\mu}}\boldsymbol{\tilde{\mu}}^{\top}+\left(\tilde{w}t\tilde{\sigma}^{2}_{r,t}+\lambda\right)\boldsymbol{I}-\sum_{i=1}^{K}\omega\sqrt{N_{i,t+1}\log\left(\frac{t}{\alpha}\right)}\boldsymbol{ee}^{\top}\right)\boldsymbol{x}$
@@
where (a) holds by Eq. 33 and Eq. 34, (b) holds since the squared root term is maximized when $N_{i,t+1}=t/K,\forall i\in[K]$. Note that $\boldsymbol{B}_{t}$ is diagonal matrix, and $\boldsymbol{C}_{t}$ is rank-1 matrix yields one eigenvalue of $\omega\sqrt{Kt\log\left(\frac{t}{\alpha}\right)}\|\boldsymbol{e}\|_{2}^{2}=-D\omega\sqrt{Kt\log\left(\frac{t}{\alpha}\right)}$ and D1D-1 eigenvalues of 0, we have
@@
Note that $\boldsymbol{A}_{t}$ and $\boldsymbol{C}_{t}$ are rank-1 matrices and $\boldsymbol{B}_{t}$ is diagonal matrix, we can trivially derive that:
  • $\boldsymbol{A}_{t}+\boldsymbol{B}_{t}$ has one eigenvalue of $\tilde{w}t(\|\boldsymbol{\tilde{\mu}}\|_{2}^{2}+\tilde{\sigma}^{2}_{r,t})+\lambda$ and D1D-1 eigenvalues of $\tilde{w}t\tilde{\sigma}^{2}_{r,t}+\lambda$.
  • $\boldsymbol{C}_{t}$ has one eigenvalue of $\omega\sqrt{Kt\log\left(\frac{t}{\alpha}\right)}\|\boldsymbol{e}\|_{2}^{2}=-D\omega\sqrt{Kt\log\left(\frac{t}{\alpha}\right)}$ and D1D-1 eigenvalues of 0.
Since $\boldsymbol{A}_{t}+\boldsymbol{B}_{t}$ and $\boldsymbol{C}_{t}$ are both symmetric, by applying Lemma 11, we have
@@ D.7 Proof of Lemma 10
$\sum_{t=M+1}^{\top}\sqrt{\frac{\log\left(\frac{t}{\alpha}\right)}{N_{a_{t},t}}}$ $\underset{(a)}{\leq}\sqrt{\log\left(\frac{T}{\alpha}\right)}\sum_{i=1}^{K}\sum_{n=N_{i,M+1}}^{N_{i,T}}\sqrt{\frac{1}{n}}$ (39)
@@ D.8 Proof of Lemma 9
To derive the upper-bound of term $\sum_{t=M+1}^{\top}\min\left(\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}\left[\boldsymbol{V}_{t-1}\right]^{-1}},\frac{1}{2\sqrt{2}}\right)$, we follow the similar techniques for analyzing the sum of instantaneous regret in OFUL (Abbasi-Yadkori, Pál, and Szepesvári 2011). Specifically, we first show that the sum of squared terms $\min\left(\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}\left[\boldsymbol{V}_{t-1}\right]^{-1}},\frac{1}{2\sqrt{2}}\right)^{2}$ is optimal up to $\mathcal{O}(\log T)$, and then extend the result to the sum of $\min\left(\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}\left[\boldsymbol{V}_{t-1}\right]^{-1}},\frac{1}{2\sqrt{2}}\right)$ using Cauchy-Schwarz inequality.
@@
$\emph{det}\left(\mathbb{E}[\boldsymbol{V}_{T}]\right)\geq\emph{det}\left(\mathbb{E}[\boldsymbol{V}_{M}]\right)\prod_{t=M+1}^{\top}\left(1+\frac{\emph{det}\left(\Sigma_{r,a_{t}}\right)}{\emph{det}\left(\mathbb{E}[\boldsymbol{V}_{t-1}]\right)}+\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t-1}]^{-1}\boldsymbol{\mu}_{a_{t}}\right).$
Please see Appendix D.8 for the detailed proof of Lemma 13.
@@
Please see Appendix D.8 for the detailed proof of Lemma 14.
@@
$\sum_{t=M+1}^{\top}\min\left(\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}\left[\boldsymbol{V}_{t-1}\right]^{-1}},\frac{1}{2\sqrt{2}}\right)^{2}$ (40)
$\qquad\qquad=\sum_{t=M+1}^{\top}\min\left(\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t}]^{-1}\boldsymbol{\mu}_{a_{t}},\frac{1}{8}\right)$
$\qquad\qquad\underset{(a)}{\leq}\sum_{t=M+1}^{\top}\frac{1}{8\log(9/8)}\log\left(1+\min\big(\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t}]^{-1}\boldsymbol{\mu}_{a_{t}},\frac{1}{8}\big)\right)$
$\qquad\qquad\leq\sum_{t=M+1}^{\top}\frac{1}{8\log(9/8)}\log\left(1+\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t}]^{-1}\boldsymbol{\mu}_{a_{t}}\right)$
@@
$\log\left(\frac{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{T}]\right)}{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{M}]\right)}\right)\geq\sum_{t=M+1}^{\top}\log\left(1+\frac{\text{det}\left(\Sigma_{r,a_{t}}\right)}{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{t-1}]\right)}+\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t-1}]^{-1}\boldsymbol{\mu}_{a_{t}}\right).$ (41)
@@
$D\log\left(1+\frac{\omega}{\lambda}(T-M)\right)\geq\log\left(\frac{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{T}]\right)}{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{M}]\right)}\right)\geq\sum_{t=M+1}^{\top}\log\left(1+\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t}]^{-1}\boldsymbol{\mu}_{a_{t}}\right).$ (42)
@@
$\sum_{t=M+1}^{\top}\min\left(\sqrt{\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t}]^{-1}\boldsymbol{\mu}_{a_{t}}},\frac{1}{2\sqrt{2}}\right)^{2}$ $\leq\frac{D}{8\log(9/8)}\log\left(1+\frac{\omega}{\lambda}(T-M)\right).$ (43)
@@
$\sum_{t=M+1}^{\top}\min\Big(\sqrt{\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t}]^{-1}\boldsymbol{\mu}_{a_{t}}},\frac{1}{2\sqrt{2}}\Big)$ $\leq\sqrt{(T-M)\sum_{t=M+1}^{\top}\min\Big(\sqrt{\boldsymbol{\mu}_{a_{t}}^{\top}\mathbb{E}[\boldsymbol{V}_{t}]^{-1}\boldsymbol{\mu}_{a_{t}}},\frac{1}{2\sqrt{2}}\Big)^{2}}$ (44)
@@
For the proof of this lemma, we follow the main idea of Determinant-Trace Inequality in OFUL (Abbasi-Yadkori, Pál, and Szepesvári 2011) (Lemma 10). Specifically, by the definition of $\boldsymbol{V}_{t}$, we have
$\log\left(\frac{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{T}]\right)}{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{M}]\right)}\right)$ $=\log\left(\text{det}\left(\frac{\mathbb{E}[\boldsymbol{V}_{M}]+\sum_{t=M+1}^{\top}\mathbb{E}[w_{t}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}]}{\mathbb{E}[\boldsymbol{V}_{M}]}\right)\right)$ (49)
$\underset{(a)}{\leq}\log\left(\text{det}\left(1+\frac{\sum_{t=M}^{\top}\mathbb{E}[w_{t}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}]}{\lambda\boldsymbol{I}}\right)\right)$
$=\log\left(\text{det}\left(1+\frac{1}{\lambda}\sum_{t=M+1}^{\top}\mathbb{E}[w_{t}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}]\right)\right),$
where (a) holds since $\text{det}(\mathbb{E}[\boldsymbol{V}_{M}])\geq\text{det}(\mathbb{E}[\boldsymbol{V}_{0}])=\lambda\boldsymbol{I}$. Let $\xi_{1},...,\xi_{D}$ denote the eigenvalues of $\sum_{t=M+1}^{\top}\mathbb{E}[w_{t}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}]$, and note:
$\sum_{d=1}^{D}\xi_{d}$ $=\text{Trace}\left(\sum_{t=M+1}^{\top}\mathbb{E}[w_{t}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}]\right)$ (50)
$=\sum_{t=M+1}^{\top}\mathbb{E}\left[\text{Trace}\left(w_{t}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}\right)\right]$
$=\sum_{t=M+1}^{\top}\mathbb{E}\left[\text{Trace}\left(\frac{\omega}{\|\boldsymbol{r}_{a_{t},t}\|_{2}^{2}}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}\right)\right]$
$=\sum_{t=M+1}^{\top}\mathbb{E}\left[\frac{\omega\|\boldsymbol{r}_{a_{t},t}\|_{2}^{2}}{\|\boldsymbol{r}_{a_{t},t}\|_{2}^{2}}\right]$
@@
$\log\left(\frac{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{T}]\right)}{\text{det}\left(\mathbb{E}[\boldsymbol{V}_{M}]\right)}\right)$ $\leq\log\left(\text{det}\left(1+\frac{1}{\lambda}\sum_{t=M+1}^{\top}\mathbb{E}[w_{t}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}]\right)\right)$